# Small-time asymptotics and the emergence of complex singularities for the KdV equation

### v1

@@ 5 Discussion
  • As an integrable pde, the KdV equation (1) is well known to be exceptional, with various special properties. Clearly, in our study, the integrability of the KdV equation plays a role in the analysis of the inner problem for $f_{0}(\xi)$, which is a change of variables away from PII, most notably in the absence of branch points but also reflected in the Stokes multipliers being explicit. On the other hand, the same general asymptotic framework we outline here for studying the emergence of complex-plane singularities should apply to other (non-integrable) dispersive wave equations, albeit with a stronger reliance on nontrivial numerical analysis of inner problems. Further, it appears that integrability does not significantly affect the application of exponential asymptotics to derive the form of the dispersive waves (5), suggesting that these techniques are likely to be versatile in this context. As such, we are motivated to expand our study to other dispersive wave equations, whether they be integrable or not.
  • An important point to emphasise is that various parts of our analysis (or the analysis for more general initial conditions with (82), for example) require that the initial condition $u_{0}(x)$ be an analytic function of $x$. If we drop that restriction, then we are unable to apply our methodologies directly to determine the small-time behaviour of the dispersive waves or the singularities of the solution in the complex plane. For example, for non-analytic initial conditions, we are not able by the above arguments to make any claim about the form of the dispersive waves for early times, or even whether the waves are exponentially small in time. Further, for non-analytic initial conditions we do not expect the early-time singularity structure to be related to the decreasing tritronquée solutions of PII, or even necessarily related to PII at all. With this in mind, a challenge would be to construct an appropriate analysis of the small-time behaviour of KdV solutions on the real line and the complex plane for non-analytic initial conditions.
@@
Acknowledgements: SWM is grateful for valuable discussions with Bernard Deconinck, Sergey Dyachenko, Tom Trogdon and Catherine Johnston, as well as to Bengt Fornberg and Andre Weideman for kindly providing a copy of the code they developed for Painlevé II [40]. SWM and CJL thank Inês Aniceto for discussions about transseries. SWM, CJL and SJC acknowledge the support of Australian Research Council Discovery Projects DP250101095, DP240101666 and DP190101190. JRK gratefully acknowledges a Royal Society Leverhulme Trust Senior Research Fellowship. SWM, CJL, JRK and SJC would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programmes ‘Complex analysis: techniques, applications and computations’ and ‘Applicable resurgent asymptotics: towards a universal theory’, where part of the work on this paper was undertaken. These programmes were supported by the EPSRC grant no. EP/R014604/1.
@@ A.1 Late-order terms in power-series expansion
We begin by writing out $u(x,t)$ as a regular perturbation in powers of $t$, as in (10), and then substituting into the KdV equation (1) to determine the correction term $u_{1}$ and subsequent terms via the recurrence relation (11). Such an asymptotic expansion is divergent due to the repeated differentiation required to compute $u_{n+1}$ from $u_{n}$. Therefore, the terms in (10) become increasingly singular in the neighbourhood of any singularity of the leading-order term $u_{0}$. For our problem, this leading-order term is simply the initial condition for the original pde.
The divergence in (10) can be characterised by analysing late-order terms $u_{n}$ as $n\rightarrow\infty$. One way to do this is to pose the factorial-over-power ansatz (12), where $\mathcal{A}$ and $\chi$ are functions to be determined and $\gamma$ is an unknown constant [16]. By differentiating (12) we find
@@
$u_{n}^{\prime\prime\prime}\sim-\frac{\mathcal{A}\chi^{\prime 3}\Gamma(2n+\gamma+3)}{\chi^{2n+\gamma+3}}+\frac{(3\mathcal{A}\chi^{\prime}\chi^{\prime\prime}+3\mathcal{A}^{\prime}\chi^{\prime 2})\Gamma(2n+\gamma+2)}{\chi^{2n+\gamma+2}},$
as $n\rightarrow\infty$, so that the recurrence relation (11) can be used to show that $\chi$ and $\mathcal{A}$ satisfy (13). The first of these equations suggests $\chi=\pm(2/3^{3/2})(x-x_{0})^{3/2}$, where recall we are assuming that the singularity $x_{0}$ is in the upper half plane. We choose the negative sign so that $\chi$ is real and positive on the Stokes line $\mathrm{arg}(x-x_{0})=-2\pi/3$. It is interesting to note that, in order to derive (13), we need only balance the first two terms of $(n+1)u_{n+1}$ with the first two terms of $u_{n}^{\prime\prime\prime}$ in the limit $n\rightarrow\infty$, which means the nonlinear terms on the right-hand side of (11) do not play a role in this part of the analysis.
Given the form of the late-order terms, (15), we can determine by $\gamma$ by a consistency check with the leading order term $u_{0}$. As we are focussing on the initial conditions with the property (3), we see that by setting $n=0$ in the denominator of (15), we have $\gamma+1/2=2$. While we do not consider other possibilities here, it is clear that other types of initial conditions would lead to different values of $\gamma$.
A.2 Switching on an exponential across Stokes lines
In this subsection, we show that the term (16) is switched on across the Stokes line $\mathrm{arg}(x-x_{0})=-2\pi/3$. A starting point is to truncate our divergent series (10) via
$u=\sum_{n=0}^{N-1}t^{n}u_{n}(x)+R_{N}(x,t).$ (84)
Here, $R_{N}$ is the remainder term which, for fixed $N$, is of $\mathcal{O}(t^{-1})$ as $t\rightarrow 0^{+}$. However, if we truncate optimally (at the smallest term), then the remainder is exponentially small. This result motivates what follows. To complicate matters, the optimal truncation point, $N=N_{\mathrm{opt}}$, depends on $x$ and $t$. For a fixed $x$, as we decrease $t$ we need more terms in our optimally truncated series. Thus, we have $N_{\mathrm{opt}}\rightarrow\infty$ as $t\rightarrow 0^{+}$, which means that all of the crucial information is in the tail of the series (10).
Substituting (84) into the KdV equation (1) gives
@@ A.2 Switching on an exponential across Stokes lines
which, after cancelling the first N1N-1 terms in each series using (11), simplifies to
@@
The ellipsis denote terms that will be smaller in size as $t\rightarrow 0^{+}$. Again, appealing to (11), our equation for the remainder becomes
$\frac{\partial R_{N}}{\partial t}+6\frac{\partial}{\partial x}\left(u_{0}R_{N}\right)+\frac{\partial^{3}R_{N}}{\partial x^{3}}\sim\frac{t^{N-1}\chi\mathcal{A}\,\Gamma(2N+\gamma+1)}{2\chi^{2N+\gamma+1}}\quad\mbox{as}\quad t\rightarrow 0^{+},$ (85)
where, for the right-hand side, we have used (12) to derive a leading order approximation for $Nt^{N-1}u_{N}$.
Note that in (85), and in what follows, the amplitude function $\mathcal{A}$ and the singulant $\chi$ are given by (14). These functions also arise as ingredients of the solution to the homogeneous part of (85). Indeed, applying a Liouville–Green method reveals that the homogenous solution, $R_{\mathrm{H}}$, behaves like
@@
To proceed, we apply a type of variation of parameters argument, and set
@@
$(\mathcal{S}_{t}+3t^{-1}\chi^{\prime 2}\mathcal{S}_{x})t^{-\gamma/2}\mathrm{e}^{-\chi/t^{1/2}}\sim\frac{t^{N-1}\chi\Gamma(2N+\gamma+1)}{2\chi^{2N+\gamma+1}}.$
@@
and the first equation in (13), we find
$\frac{\mathrm{d}\mathcal{S}}{\mathrm{d}\bar{\chi}}\sim\frac{\Gamma(2N+\gamma+1)}{2\bar{\chi}^{2N+\gamma+1}}\,\mathrm{e}^{\bar{\chi}}\quad\mbox{as}\quad|\bar{\chi}|\rightarrow\infty.$ (86)
@@
In the above analysis, we wish to take $N=N_{\mathrm{opt}}$, where $N_{\mathrm{opt}}$ comes from truncating (84) optimally. We follow a heuristic which says optimal truncation occurs at the least term or, in other words, when
@@
Using (12), we find that $N_{\mathrm{opt}}\sim|\chi|/2t^{1/2}$, so we write
@@
Before we substitute in our optimal value of $N$ into (86), we make a change of variables $\bar{\chi}=\bar{\rho}\mathrm{e}^{\mathrm{i}\vartheta}$. Since $N_{\mathrm{opt}}$ depends on $|\bar{\chi}|=\bar{\rho}$ and not $\vartheta$, then we rewrite our $\bar{\chi}$ derivative as
@@
Using Stirling’s formula, our equation (86) for the Stokes multiplier becomes
@@
The argument $1-\mathrm{i}\vartheta+\mathrm{e}^{\mathrm{i}\vartheta}$ has a negative real part, except when $\vartheta=0$. Thus, we see that the rate of change of the Stokes multiplier, $\mathrm{d}\mathcal{S}/\mathrm{d}\vartheta$, is exponentially small for $\vartheta=\mathcal{O}(1)$, $\bar{\rho}\gg 1$, except at $\vartheta=0$ where it is of $\mathcal{O}(\bar{\rho}^{1/2})$, suggesting a boundary layer near $\vartheta=0$.
@@
In terms of the variable $\bar{\theta}$, the jump in $\mathcal{S}$ across $\bar{\theta}=0$ is smooth, approximated by an error function [8]. However, from the perspective of the original spatial variable $x$, the transition appears to be sharp and the observation is that we pick up a contribution $\mathcal{S}]_{-}^{+}$ multiplied by $\mathcal{A}\,t^{-\gamma/2}\mathrm{e}^{-\chi/t^{1/2}}$ as we cross the Stokes line $\bar{\theta}=0$. Now, the function $\chi=-(2/3^{3/2})(x-x_{0})^{3/2}$ is real and positive when $\sin\left(\frac{3}{2}\mathrm{arg}(x-x_{0})\right)=0$ and $\cos\left(\frac{3}{2}\mathrm{arg}(x-x_{0})\right)<0$. Thus, we have shown that (16) switches on as we cross the Stokes line $\mathrm{arg}(x-x_{0})=-2\pi/3$.
A.3 Matching into the inner region to determine $\Lambda$
Finally, we need to determine the Stokes constant $\Lambda$, which we do by matching into the inner region near $x=x_{0}$. To begin, we employ the inner variable $\xi$, defined in (23), and construct the inner limit of the outer expansion using
$t^{n}u_{n}\sim\frac{\Lambda t^{n}(-3^{3/2}/2)^{2n+3/2}\Gamma(2n+{\textstyle\frac{3}{2}})}{(x-x_{0})^{3n+2}}=\Lambda\left(\frac{-3^{3/2}}{2}\right)^{3/2}\left(\frac{3}{2}\right)^{2n}\frac{\Gamma(2n+{\textstyle\frac{3}{2}})}{(3t)^{2/3}\xi^{3n+2}}$ (87)

### v2

@@ 5 Discussion
Of course, there are many other classes of initial conditions that do not have singularities of the type (5.2), and these are also of interest for future work. For example, entire initial conditions do not have any singularities in the plane at all and, thus, for such cases, the singularities of the time-dependent solution must be born at infinity. Say we had a Gaussian initial condition $u_{0}=A\,\mathrm{e}^{-x^{2}}$, then from (2.2) we have $u_{1}=12A^{2}x\,\mathrm{e}^{-2x^{2}}-4Ax(3-2x^{2})\,\mathrm{e}^{-x^{2}}$. As we move up into the upper half of the complex-$x$ plane, both terms of $u_{1}$ grow exponentially, but the first dominates the second. Indeed, the first term in $u_{1}$ grows faster than $u_{0}$ does in this direction. Thus, the expansion (2.1) breaks down where $A\,\mathrm{e}^{-x^{2}}=\mathcal{O}(12A^{2}tx\,\mathrm{e}^{-2x^{2}})$, which suggests an inner problem with scalings
$u=-\frac{\mathrm{i}}{12t\ln^{1/2}(1/t)}U(X,t),\quad x=\mathrm{i}\ln^{1/2}(1/t)+\frac{-\frac{1}{4}\mathrm{i}\ln(\ln(1/t))-\frac{1}{2}\mathrm{i}\ln(12A)+(n+\frac{1}{4})\pi+X}{\ln^{1/2}(1/t)},$
where $X=\mathcal{O}(1)$ and $n$ is an integer whose magnitude is of $\mathcal{O}(1)$. To leading order, we write $U\sim U_{0}(X)$ as $t\rightarrow 0^{+}$, where
$U_{0}+{\textstyle\frac{1}{2}}\mathrm{i}\frac{\mathrm{d}U_{0}}{\mathrm{d}X}-{\textstyle\frac{1}{2}}\mathrm{i}U_{0}\frac{\mathrm{d}U_{0}}{\mathrm{d}X}=0.$ (5.4)
As the third-order dispersive term in KdV does not appear in (5.4), the resulting singularities of $U_{0}$ will be of the wrong type (they will be square-root-type branch points instead of double poles). Thus, there will need to be a further rescaling near these singularities to establish the appropriate balance between advection and dispersion. This analysis will complicate the description of the early-time pole dynamics as well as the exponential asymptotics. Other entire initial conditions are expected to involve similar challenges.
@@
As an integrable pde, the KdV equation (1.1) is well known to be exceptional, with various special properties. Clearly, in our study, the integrability of the KdV equation plays a role in the analysis of the inner problem for $f_{0}(\xi)$, which is a change of variables away from PII, most notably in the absence of branch points but also reflected in the Stokes multipliers being explicit. On the other hand, the same general asymptotic framework we outline here for studying the emergence of complex-plane singularities should apply to other (non-integrable) dispersive wave equations, albeit with a stronger reliance on nontrivial numerical analysis of inner problems. Further, it appears that integrability does not significantly affect the application of exponential asymptotics to derive the form of the dispersive waves (1.5), suggesting that these techniques are likely to be versatile in this context. As such, we are motivated to expand our study to other dispersive wave equations, whether they be integrable or not.
An important point to emphasise is that various parts of our analysis (or the analysis for more general initial conditions with (5.2), for example) require that the initial condition $u_{0}(x)$ be an analytic function of $x$. If we drop that restriction, then we are unable to apply our methodologies directly to determine the small-time behaviour of the dispersive waves or the singularities of the solution in the complex plane. For example, for non-analytic initial conditions, we are not able by the above arguments to make any claim about the form of the dispersive waves for early times, or even whether the waves are exponentially small in time. Further, for non-analytic initial conditions we do not expect the early-time singularity structure to be related to the decreasing tritronquée solutions of PII, or even necessarily related to PII at all. With this in mind, a challenge would be to construct an appropriate analysis of the small-time behaviour of KdV solutions on the real line and the complex plane for non-analytic initial conditions.
@@
{Backmatter}
Acknowledgments
SWM is grateful for valuable discussions with Bernard Deconinck, Sergey Dyachenko, Tom Trogdon and Catherine Johnston, as well as to Bengt Fornberg and Andre Weideman for kindly providing a copy of the code they developed for Painlevé II [46]. SWM and CJL thank Inês Aniceto for discussions about transseries. SWM, CJL, JRK and SJC would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programmes ‘Complex analysis: techniques, applications and computations’ and ‘Applicable resurgent asymptotics: towards a universal theory’, where part of the work on this paper was undertaken. These programmes were supported by the EPSRC grant no. EP/R014604/1. All authors thank the anonymous referees for their constructive comments and suggestions. For the purpose of open access, the authors have applied a CC BY public copyright licence to any author accepted manuscript arising from this submission.
Funding Statement
SWM, CJL and SJC acknowledge the support of Australian Research Council Discovery Projects DP250101095, DP240101666 and DP190101190. JRK gratefully acknowledges a Royal Society Leverhulme Trust Senior Research Fellowship.
Competing Interests
None declared.
Data Availability Statement
The datasets used in this study are available from the authors upon request.
Ethical Standards
The research meets all ethical guidelines, including adherence to the legal requirements of the study country.
Author Contributions
Conceptualization: SWM, CJL, JRK, SJC. Methodology: SWM, CJL, DJV, JZ, JRK, SJC. Data curation: SWM, CJL, DJV, JZ, JRK, SJC. Data visualisation: CJL, DJV, JZ. Writing original draft: SWM. All authors approved the final submitted draft.
@@ A.1 Late-order terms in power-series expansion
We begin by writing out $u(x,t)$ as a regular perturbation in powers of $t$, as in (2.4), and then substituting into the KdV equation (1.1) to determine the correction term $u_{1}$ and subsequent terms via the recurrence relation (2.5). Such an asymptotic expansion is divergent due to the repeated differentiation required to compute $u_{n+1}$ from $u_{n}$. Therefore, the terms in (2.4) become increasingly singular in the neighbourhood of any singularity of the leading-order term $u_{0}$. For our problem, this leading-order term is simply the initial condition for the original pde.
The divergence in (2.4) can be characterised by analysing late-order terms $u_{n}$ as $n\rightarrow\infty$. One way to do this is to pose the factorial-over-power ansatz (2.6), where $\mathcal{A}$ and $\chi$ are functions to be determined and $\gamma$ is an unknown constant [20]. By differentiating (2.6) we find
@@
$u_{n}^{\prime\prime\prime}\sim-\frac{\mathcal{A}(\chi^{\prime})^{3}\Gamma(2n+\gamma+3)}{\chi^{2n+\gamma+3}}+\frac{(3\mathcal{A}\chi^{\prime}\chi^{\prime\prime}+3\mathcal{A}^{\prime}(\chi^{\prime})^{2})\Gamma(2n+\gamma+2)}{\chi^{2n+\gamma+2}},$
as $n\rightarrow\infty$, so that the recurrence relation (2.5) can be used to show that $\chi$ and $\mathcal{A}$ satisfy (2.7). The first of these equations suggests $\chi=\pm(2/3^{3/2})(x-x_{0})^{3/2}$, where recall we are assuming that the singularity $x_{0}$ is in the upper half plane. We choose the negative sign so that $\chi$ is real and positive on the Stokes line $\mathrm{arg}(x-x_{0})=-2\pi/3$. It is interesting to note that, in order to derive (2.7), we need only balance the first two terms of $(n+1)u_{n+1}$ with the first two terms of $u_{n}^{\prime\prime\prime}$ in the limit $n\rightarrow\infty$, which means the nonlinear terms on the right-hand side of (2.5) do not play a role in this part of the analysis.
Given the form of the late-order terms, (2.9), we can determine by $\gamma$ by a consistency check with the leading order term $u_{0}$. As we are focussing on the initial conditions with the property (1.3), we see that by setting $n=0$ in the denominator of (2.9), we have $\gamma+1/2=2$. While we do not consider other possibilities here, it is clear that other types of initial conditions would lead to different values of $\gamma$.
A.2 Switching on exponential across Stokes lines
In this subsection, we show that the term (2.10) is switched on across the Stokes line $\mathrm{arg}(x-x_{0})=-2\pi/3$. A starting point is to truncate out divergent series (2.4) so that
$u=\sum_{n=0}^{N-1}t^{n}u_{n}(x)+R_{N}(x,t).$ (A.1)
Here, $R_{N}$ is the remainder term which, for fixed $N$, is of $\mathcal{O}(t^{-1})$ as $t\rightarrow 0^{+}$. However, if we truncate optimally (at the smallest term), then the remainder is exponentially small. This result motivates what follows. To complicate matters, the optimal truncation point, $N=N_{\mathrm{opt}}$, depends on $x$ and $t$. For a fixed $x$, as we decrease $t$ we need more terms in our optimally truncated series. Thus, we have $N_{\mathrm{opt}}\rightarrow\infty$ as $t\rightarrow 0^{+}$, which means that all of the crucial information is in the tail of the series (2.4).
Substituting (A.1) into the KdV equation (1.1) gives
@@ A.2 Switching on exponential across Stokes lines
which, after cancelling the first N1N-1 terms in each series using (2.5), simplifies to
@@
The ellipsis denote terms that will be smaller in size as $t\rightarrow 0^{+}$. Again, appealing to (2.5), our equation for the remainder becomes
$\frac{\partial R_{N}}{\partial t}+6\frac{\partial}{\partial x}\left(u_{0}R_{N}\right)+\frac{\partial^{3}R_{N}}{\partial x^{3}}\sim\frac{t^{N-1}\chi\mathcal{A}\,\Gamma(2N+\gamma+1)}{2\chi^{2N+\gamma+1}}\quad\mbox{as}\quad t\rightarrow 0^{+},$ (A.2)
where, for the right-hand side, we have used (2.6) to derive a leading order approximation for $Nt^{N-1}u_{N}$.
Note that in (A.2), and in what follows, the amplitude function $\mathcal{A}$ and the singulant $\chi$ are given by (2.8). These functions also arise as ingredients of the solution to the homogeneous part of (A.2). Indeed, applying a WKB method reveals that the homogenous solution, $R_{\mathrm{H}}$, behaves like
@@
To proceed, we apply a type of variation of parameters argument, and let
@@
$(\mathcal{S}_{t}+3t^{-1}(\chi^{\prime})^{2}\mathcal{S}_{x})t^{-\gamma/2}\mathrm{e}^{-\chi/t^{1/2}}\sim\frac{t^{N-1}\chi\,\Gamma(2N+\gamma+1)}{2\chi^{2N+\gamma+1}}.$
@@
and the first equation in (2.7), we find
$\frac{\mathrm{d}\mathcal{S}}{\mathrm{d}\bar{\chi}}\sim\frac{\Gamma(2N+\gamma+1)}{2\bar{\chi}^{2N+\gamma+1}}\,\mathrm{e}^{\bar{\chi}}\quad\mbox{as}\quad|\bar{\chi}|\rightarrow\infty.$ (A.3)
@@
In the above analysis, we wish to take $N=N_{\mathrm{opt}}$, where $N_{\mathrm{opt}}$ comes from truncating (A.1) optimally. We follow a heuristic which says optimal truncation occurs at the least term or, in other words, when
@@
Using (2.6), we find that $N_{\mathrm{opt}}\sim|\chi|/2t^{1/2}$, so we write
@@
Before we substitute in our optimal value of $N$ into (A.3), we make a change of variables $\bar{\chi}=\bar{\rho}\mathrm{e}^{\mathrm{i}\vartheta}$. Since $N_{\mathrm{opt}}$ depends on $|\bar{\chi}|=\bar{\rho}$ and not $\vartheta$, then we rewrite our $\bar{\chi}$ derivative as
@@
Using Stirling’s formula, our equation (A.3) for the Stokes multiplier becomes
@@
The argument $1-\mathrm{i}\vartheta+\mathrm{e}^{\mathrm{i}\vartheta}$ has a negative real part, except when $\vartheta=0$. Thus, we see that the rate of change of the Stokes multiplier, $\mathrm{d}\mathcal{S}/\mathrm{d}\vartheta$, is exponentially small for $\bar{\rho}\gg 1$, except at $\vartheta=0$ where it is of $\mathcal{O}(\bar{\rho}^{1/2})$, suggesting a boundary layer near $\vartheta=0$.
@@
In terms of the variable $\bar{\theta}$, the jump in $\mathcal{S}$ across $\bar{\theta}=0$ is smooth, approximated by an error function [9]. However, from the perspective of the original spatial variable $x$, the transition appears to be sharp and the observation is that we pick up a contribution $\mathcal{S}]_{-}^{+}$ multiplied by $\mathcal{A}\,t^{-\gamma/2}\mathrm{e}^{-\chi/t^{1/2}}$ as we cross the Stokes line $\bar{\theta}=0$. Now, the function $\chi=-(2/3^{3/2})(x-x_{0})^{3/2}$ is real and positive when $\sin(\frac{3}{2}\mathrm{arg}(x-x_{0}))=0$ and $\cos(\frac{3}{2}\mathrm{arg}(x-x_{0}))<0$. Thus, we have shown that (2.10) switches on as we cross the Stokes line $\mathrm{arg}(x-x_{0})=-2\pi/3$.
A.3 Matching in inner region to determine $\Lambda$
Finally, we need to determine the Stokes constant $\Lambda$, which we do by matching into the inner region near $x=x_{0}$. To begin, we employ the inner variable $\xi$, defined in (3.1), and construct the inner limit of the outer expansion using
$t^{n}u_{n}\sim\frac{\Lambda t^{n}(-3^{3/2}/2)^{2n+3/2}\Gamma(2n+{\textstyle\frac{3}{2}})}{(x-x_{0})^{3n+2}}=\Lambda\left(\frac{-3^{3/2}}{2}\right)^{3/2}\left(\frac{3}{2}\right)^{2n}\frac{\Gamma(2n+{\textstyle\frac{3}{2}})}{(3t)^{2/3}\xi^{3n+2}}$ (A.4)